OpenStudy (anonymous):

Trying to solve the following: Fran works due north of home. Her husband Ray works due East. They leave at the same time. By the time Fran is 7 miles from home, the distance between them is one mile more than Ray's distance from home. How far from home is Ray?

OpenStudy (dumbcow):

 use pythagorean thm to solve for x

OpenStudy (anonymous):

so we add x+x+1+7

OpenStudy (anonymous):

2x+8

OpenStudy (anonymous):

x=4?

OpenStudy (dumbcow):

where is the "=" sign ?? haha you cant just say 2x+8 , then x = 4 anyway pythagorean theorem for right triangles is: \[a^2 + b^2 = c^2\] \[7^2 + x^2 = (x+1)^2\]

OpenStudy (anonymous):

I'm lost

OpenStudy (dumbcow):

this is pretty basic, you should have already learned about pythagorean thm 

OpenStudy (anonymous):

not in this type of problem